Calculate the temperature at which the resistance of a conductor becomes 20% more than its resistance at 27 degree Celsius. The Value of temperature Coefficient of resistance of the conductor is 2×10-4  Per Kelvin.

Grade:12

1 Answers

Athul Joseph
5 years ago
R`=R(1+@(Tf-Ti)).R`=1.2RTi=300 K@=0.0002On solving,1.2R=R(1+0.0002(Tf-300))0.2=0.0002 (Tf-300)Tf=1000+300We get Tf=1027 °C or 1300K